# Procedural Mistake Detection via Action Effect Modeling

### v1

@@ Abstract
Mistake detection in procedural tasks is essential for building intelligent systems that support learning and task execution. Existing approaches primarily analyze how an action is performed, while overlooking what it produces, i.e., the action effect. Yet many errors manifest not in the execution itself but in the resulting outcome, such as an unintended object state or incorrect spatial arrangement. To address this gap, we propose Action Effect Modeling (AEM), a unified framework that jointly captures action execution and its outcomes through a probabilistic formulation. AEM first identifies the outcome of an action by selecting the most informative effect frame based on semantic relevance and visual quality. It then extracts complementary cues from visual grounding and symbolic scene graphs, aligning them in a shared latent space to form robust effect-aware representations. To detect mistakes, we further design a prompt-based detector that incorporates task-specific prompts and aligns each action segment with its intended execution semantics. Our approach achieves state-of-the-art performance on the EgoPER and CaptainCook4D benchmarks under the challenging one-class classification (OCC) setting. These results demonstrate that modeling both execution and outcome yields more reliable mistake detection, and highlight the potential of effect-aware representations to benefit a broader range of downstream applications.
@@ 1 Introduction
Existing mistake detection approaches focus primarily on modeling the execution process, assessing correctness by analyzing motion patterns and action sequences. Previous works have explored such methods in structured tasks, such as assembly (sener2022assembly101) and cooking (wang2023holoassist; peddi2023captaincook4d), often relying on frame-level classification or multimodal features. More recently, lee2024error introduced prototype learning to evaluate execution similarity across instances, while huang2025modeling utilized task graphs to predict all possible next actions and compared them with actual executions in the latent space. However, a shared limitation of these methods is the assumption that mistakes can be identified solely from the execution process, without verifying whether the final outcome aligns with the intended result.
@@
To learn robust and interpretable effect-aware representations, AEM leverages multimodal supervision from vision-language models (hurst2024gpt) through dual branches. The visual branch encodes grounded object features and positions, while a textual branch constructs symbolic scene graphs to extract attributes and relational descriptions. During training, these supervision signals are aligned in a shared latent space via contrastive objectives, allowing the model to distill structured effect knowledge into a compact representation. At test time, the learned effect representation is fused with the action segment features, providing enriched context for downstream mistake detection.
While action effects provide valuable cues, effective mistake detection also requires modeling the execution itself. To this end, we develop a prompt-based detector that evaluates whether an action segment aligns with its execution process within the task context. Rather than relying on frame-level classifiers, our detector encodes each action using a learnable representation and aligns it with a task-specific textual prompt in a contrastive manner. By modeling both the temporal dynamics and the resulting effect of an action, the model captures rich semantic cues to distinguish subtle execution errors as well as outcome discrepancies. Our approach is evaluated on two challenging egocentric datasets, EgoPER (lee2024error) and CaptainCook4D (peddi2023captaincook4d), achieving state-of-the-art performance. Extensive ablations further demonstrate the necessity and effectiveness of action effect modeling in improving mistake detection.
@@
Figure 1: Examples of two types of mistakes: object positional and state mistakes, in cooking scenario from CaptainCook4D dataset (peddi2023captaincook4d).
@@ 2 Related Work
Mistake Detection consists of two main branches: online detection (jang2019epic; ding2023every; narasimhan2023learning; ashutosh2024video; flaborea2024prego; plini2024ti; soran2015generating; jang2023multimodal) and offline detection (sener2022assembly101; wang2023holoassist; haneji2024egooops; peddi2023captaincook4d; schoonbeek2024industreal; ghoddoosian2023weakly; lee2024error; mazzamuto2024eyes; storks2024explainable; huang2025modeling). Online mistake detection focuses on detecting sequential mistakes in real time, while offline detection aims to recognize incorrect action execution. This work primarily addresses offline mistake detection.
Zooming into the offline detection, early approaches relied on training binary classifiers with visual or multimodal inputs (e.g., text, audio, depth) to learn implicit action representations, without explicitly modeling the underlying action dynamics (sener2022assembly101; haneji2024egooops; peddi2023captaincook4d). More recently, lee2024error introduced a prototype-based method that explicitly models actions by clustering features during training to learn representative action embeddings. huang2025modeling utilized task graphs to predict possible next steps and reconstructed action feature in the latent space. Both methods realized error detection through similarity comparison between predicted and input action features at inference time.
@@
Video Anomaly Detection (wu2019deep; gong2019memorizing; pu2022audio; wu2022self; al2024coarse; pu2024learning) is closely related to mistake detection, where models are typically trained on normal videos and identify anomalies as statistical deviations, without requiring prior knowledge of specific abnormal events. However, mistake detection in procedural videos is inherently goal-oriented, as it requires an understanding of structured workflows to assess whether an action is executed correctly. In addition, VAD is predominantly used in surveillance systems, where the camera view remains fixed and the dynamics of the scene is relatively stable. In contrast, procedural videos, especially in egocentric views, involve more fine-grained human activities, explicit scene transitions, and dynamic human motion, making action-effect modeling more challenging but crucial for building life assistive system.
@@ 3 Methodology
3.1 Problem Formulation.
Given a procedural video corresponding to a predefined task, our goal is to determine whether each segmented action is executed correctly or contains a mistake. We operate on non-overlapping segments, where each segment is denoted as a tuple $s=(t_{s},t_{e},a,y)$, with $t_{s}$ and $t_{e}$ denoting the start and end timestamps, $a$ the action label, and $y\in\{0,1\}$ the binary mistake label ($y=1$ indicates a mistake). Following the one-class classification (OCC) setting (lee2024error; huang2025modeling), the training set includes only normal actions ($y=0$), while the test set contains both correct and erroneous segments.
@@
This formulation provides a structured view of mistake detection pipeline: frame sampling, effect-aware learning, and mistake classification, bridging the gap between how an action is performed and what outcomes it produces. Next, we will build a framework following this formulation.
3.2 Framework Overview.
@@ 3.3 Action Segmentation
Following lee2024error and huang2025modeling, we adopt ActionFormer (zhang2022actionformer) as the action segmentation backbone, which outputs a feature pyramid with multi-scale temporal resolutions. Inspired by prior work (yang2024dyfadet; meng2020ar), we design a dynamic fusion module to refine temporal representations by adaptively aggregating the multi-scale features from the backbone, while preserving the hierarchical pyramid structure for segment boundary regression. This design aims to provide more discriminative segment features $\mathbf{X}$ for subsequent action effect modeling and mistake detection. Please refer to Appendix LABEL:section:a1 for details about this module.
@@ 3.4 Action Effect Modeling
Effect Frame Sampling. To identify the frame that best captures the outcome of a given action, we rank candidate frames within each action segment based on semantic relevance and visual clarity. For semantic relevance, motivated by niu2024schema, we use GPT-4o (hurst2024gpt) to generate textual descriptions of the anticipated post-action states. Then we compute the frame-wise similarity between the segment’s visual feature $\mathbf{X}$ and the description embeddings obtained from a pre-trained text encoder, followed by softmax normalization. For visual clarity, we apply the Laplacian operator to estimate the sharpness of each frame within the segment and normalize the scores between the $5th$ and $95th$ percentiles. The final ranking score combines both semantic relevance and visual clarity, and the top-ranked frame is selected as the effect frame.
@@
In the visual branch, we use Grounding DINO (liu2024grounding) to detect relevant objects, then we extract object state and relation features $\mathbf{v}_{s},\mathbf{v}_{r}$ by concatenating and projecting embeddings:
@@
Although visual grounding provides object-centric features that capture appearance and spatial cues from the effect frame, it lacks high-level semantic abstraction and relational reasoning. To address this, we introduce a textual scene graph branch to model the symbolic structure of the action effect. Given an effect frame and objects detected by Grounding DINO, we leverage the recognition and reasoning capabilities of a Multimodal Large Language Model (MLLM) by prompting GPT-4o to generate a scene graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ contains three types of nodes: object, relation, and attribute, and $\mathcal{E}$ represents the set of directed edges. We encode the textual labels of all nodes in $\mathcal{V}$ using a pre-trained text encoder to obtain node embeddings, which are then fed into a Graph Neural Network (GNN) (velivckovic2017graph) to compute context-aware features $\mathbf{t}_{v}$ for each node.
@@ 3.5 Mistake Detection
Following previous work (lee2024error; huang2025modeling), we adopt an one-class classification (OCC) setting for mistake detection, which learns canonical action patterns from correct samples only. Given the effect-enhanced segment features, we first perform an average-pooling over the enhanced effect-aware segment feature $\widetilde{\mathbf{X}}$ across temporal dimension to obtain the action embedding $\mathbf{x}_{a}$. Instead of learning a binary classifier, we introduce a prompt-based detector to increase task-specific discriminability. For each action label $a$, we construct a template-based textual prompt $\mathcal{P}_{a}$ (e.g., “An image showing [ACTION] for [TASK]”), and obtain its representation by:
@@ 3.6 Training Objective
Finally, we jointly optimize three objectives to train the model in an end-to-end fashion: the segmentation loss $\mathcal{L}^{\text{seg}}$ following zhang2022actionformer, the action-effect modeling loss $\mathcal{L}^{\text{eff}}$ and $\mathcal{L}^{\text{CL}}$, and the mistake detection loss $\mathcal{L}^{\text{det}}$. The overall training objective is defined as:
@@ 4 Experiments
Table 1: Mistake detection results on the EgoPER dataset (lee2024error). AUC and EDA are in percentage (%). Best results are bold, and second-best are underlined.
@@
Table 2: Mistake detection results on the CaptainCook4D dataset (peddi2023captaincook4d).
@@ 4.1 Experimental Setup
Mistake Datasets. We evaluate our method on two egocentric video datasets: EgoPER (lee2024error) and CaptainCook4D (huang2025modeling). The EgoPER dataset consists of 385 egocentric cooking videos spanning 5 recipe categories, with a total duration of 28 hours. Among these, 213 videos contain only correct actions, while the remaining 178 include various execution mistakes. Following lee2024error, we use 80% of the normal videos for training, 10% for validation, and the remaining 10% together with all erroneous videos for testing. The CaptainCook4D dataset contains 384 videos covering 24 recipes, with a total length of 94.5 hours. Among all recordings, 164 are labeled as normal, while the remaining 220 videos include incorrect action steps. Following huang2025modeling, we use all normal recordings for training and the erroneous ones for testing.
Evaluation Metrics. We follow prior work (lee2024error; huang2025modeling) by using Area Under the Curve (AUC) and Error Detection Accuracy (EDA) to jointly measure model’s ability to distinguish correct and erroneous actions within each video. AUC assesses frame-level discrimination between errors and non-errors, whereas EDA measures segment-level detection accuracy. Both metrics are computed by aggregating detection results across a densely sampled threshold range from 0 to 1, rather than relying on a single fixed threshold, enabling a more comprehensive evaluation of the model’s behavior across the full spectrum of thresholds.
Implementation Details. We use off-the-shelf EVA-02 CLIP (eva02) for all the image and text encoding. For the EgoPER dataset, we follow lee2024error to use additional features based on Active Object Detection (AOD). For effect frame sampling, the final score is an average of the semantic relevance and visual quality scores. Following lee2024error and huang2025modeling, we independently train and test the model in each recipe and report the average of all recipes as the final performance. Both the effect-frame sampling and multimodal knowledge extraction are performed during the data pre-processing stage to avoid additional training overhead. All experiments are conducted using an NVIDIA RTX 6000 Ada GPU (48GB memory).
@@ 4.2 Comparison with State-of-the-art
We compare our methods with video anomaly detection methods HF2-VAD (liu2021hybrid), HF2-VAD + SSPCAB (ristea2022self) and S3R (wu2022self), as well as mistake detection methods EgoPED (lee2024error) and AMNAR (huang2025modeling). As shown in Table 1, on the EgoPER dataset, our method outperforms in both metrics on most of the five tasks, with an average improvement of 5.3% on AUC and 2.3% on EDA. The results in Table 2 further demonstrate its effectiveness on the CaptainCook4D dataset, surpassing AMNAR by 2.7% in Precision and 2.3% in AUC. On both datasets, our EDA is occasionally lower than that of AMNAR. This gap is likely due to AMNAR’s use of a dynamic programming block to construct task graphs that mitigate noise in action-segment labels. Since incorporating task graphs lies beyond the scope of our work, we leave it for future exploration.
Table 3: Ablation studies (a–d), action segmentation results (e), and performance comparison with different scene graph generation models (f) on EgoPER dataset (lee2024error).

### v2

@@ Abstract
Mistake detection in procedural tasks is essential for building intelligent systems that support learning and task execution. Existing approaches primarily analyze how an action is performed, while overlooking what it produces, i.e., the action effect. Yet many errors manifest not in the execution itself but in the resulting outcome, such as an unintended object state or incorrect spatial arrangement. To address this gap, we propose Action Effect Modeling (AEM), a unified framework that jointly captures action execution and its outcomes through a probabilistic formulation. AEM first identifies the outcome of an action by selecting the most informative effect frame based on semantic relevance and visual quality. It then extracts complementary cues from visual grounding and symbolic scene graphs, aligning them in a shared latent space to form robust effect-aware representations. To detect mistakes, we further design a prompt-based detector that incorporates task-specific prompts and aligns each action segment with its intended execution semantics. Our approach achieves state-of-the-art performance on the EgoPER and CaptainCook4D benchmarks under the challenging one-class classification (OCC) setting. These results demonstrate that modeling both execution and outcome yields more reliable mistake detection, and highlight the potential of effect-aware representations to benefit a broader range of downstream applications.  [^1] Project page: https://wenliangguo.github.io/Mistake_Detection
@@ 1 Introduction
Existing mistake detection approaches focus primarily on modeling the execution process, assessing correctness by analyzing motion patterns and action sequences. Previous works have explored such methods in structured tasks, such as assembly (Sener et al., 2022) and cooking (Wang et al., 2023; Peddi et al., 2023), often relying on frame-level classification or multimodal features. More recently, Lee et al. (2024) introduced prototype learning to evaluate execution similarity across instances, while Huang et al. (2025a) utilized task graphs to predict all possible next actions and compared them with actual executions in the latent space. However, a shared limitation of these methods is the assumption that mistakes can be identified solely from the execution process, without verifying whether the final outcome aligns with the intended result.
@@
To learn robust and interpretable effect-aware representations, AEM leverages multimodal supervision from vision-language models (Hurst et al., 2024) through dual branches. The visual branch encodes grounded object features and positions, while a textual branch constructs symbolic scene graphs to extract attributes and relational descriptions. During training, these supervision signals are aligned in a shared latent space via contrastive objectives, allowing the model to distill structured effect knowledge into a compact representation. At test time, the learned effect representation is fused with the action segment features, providing enriched context for downstream mistake detection.
While action effects provide valuable cues, effective mistake detection also requires modeling the execution itself. To this end, we develop a prompt-based detector that evaluates whether an action segment aligns with its execution process within the task context. Rather than relying on frame-level classifiers, our detector encodes each action using a learnable representation and aligns it with a task-specific textual prompt in a contrastive manner. By modeling both the temporal dynamics and the resulting effect of an action, the model captures rich semantic cues to distinguish subtle execution errors as well as outcome discrepancies. Our approach is evaluated on two challenging egocentric datasets, EgoPER (Lee et al., 2024) and CaptainCook4D (Peddi et al., 2023), achieving state-of-the-art performance. Extensive ablations further demonstrate the necessity and effectiveness of action effect modeling in improving mistake detection.
@@
Figure 1: Examples of two types of mistakes: object positional and state mistakes, in cooking scenario from CaptainCook4D dataset (Peddi et al., 2023).
@@ 2 Related Work
Mistake Detection consists of two main branches: online detection (Jang et al., 2019; Ding et al., 2023; Narasimhan et al., 2023; Ashutosh et al., 2024; Flaborea et al., 2024; Plini et al., 2024; Soran et al., 2015; Jang et al., 2023) and offline detection (Sener et al., 2022; Wang et al., 2023; Haneji et al., 2024; Peddi et al., 2023; Schoonbeek et al., 2024; Ghoddoosian et al., 2023; Lee et al., 2024; Mazzamuto et al., 2024; Storks et al., 2024; Huang et al., 2025a). Online mistake detection focuses on detecting sequential mistakes in real time, while offline detection aims to recognize incorrect action execution. This work primarily addresses offline mistake detection.
Zooming into offline detection, early approaches relied on training binary classifiers with visual or multimodal input (e.g., text, audio, depth) to learn implicit action representations, without explicitly modeling the underlying action dynamics (Sener et al., 2022; Haneji et al., 2024; Peddi et al., 2023). More recently, Lee et al. (2024) introduced a prototype-based method that explicitly models actions by clustering features during training to learn representative action embeddings. Huang et al. (2025a) utilized task graphs to predict possible next steps and reconstructed action feature in the latent space. Both methods realized error detection through comparison of similarity between the predicted and input action features at inference time.
@@
Video Anomaly Detection (Wu et al., 2019; Gong et al., 2019; Pu and Wu, 2022; Wu et al., 2022; Al-Lahham et al., 2024; Pu et al., 2024) is closely related to mistake detection, where models are typically trained on normal videos and identify anomalies as statistical deviations, without requiring prior knowledge of specific abnormal events. However, mistake detection in procedural videos is inherently goal-oriented, as it requires an understanding of structured workflows to assess whether an action is executed correctly. In addition, VAD is predominantly used in surveillance systems, where the camera view remains fixed and the dynamics of the scene is relatively stable. In contrast, procedural videos, especially in egocentric views, involve more fine-grained human activities, explicit scene transitions, and dynamic human motion, making action-effect modeling more challenging but crucial for building life assistive system.
@@ 3 Methodology
3.1 Problem Formulation
Given a procedural video corresponding to a predefined task, our goal is to determine whether each segmented action is executed correctly or contains a mistake. We operate on non-overlapping segments, where each segment is denoted as a tuple $s=(t_{s},t_{e},a,y)$, with $t_{s}$ and $t_{e}$ denoting the start and end timestamps, $a$ the action label, and $y\in\{0,1\}$ the binary mistake label ($y=1$ indicates a mistake). Following the one-class classification (OCC) setting (Lee et al., 2024; Huang et al., 2025a), the training set includes only normal actions ($y=0$), while the test set contains both correct and erroneous segments.
@@ 3.1 Problem Formulation
This formulation provides a structured view of the mistake detection pipeline: frame sampling, effect-aware learning, and mistake classification, bridging the gap between how an action is performed and what outcomes it produces. Next, we will build a framework following this formulation.
3.2 Framework Overview
@@ 3.3 Action Segmentation
Following Lee et al. (2024) and Huang et al. (2025a), we adopt ActionFormer (Zhang et al., 2022) as the action segmentation backbone, which outputs a feature pyramid with multi-scale temporal resolutions. Inspired by prior work (Yang et al., 2024; Meng et al., 2020), we design a dynamic fusion module to refine temporal representations by adaptively aggregating the multi-scale features from the backbone, while preserving the hierarchical pyramid structure for segment boundary regression. This design aims to provide more discriminative segment features $\mathbf{X}$ for subsequent action effect modeling and mistake detection. Please refer to Appendix A for details about this module.
@@ 3.4 Action Effect Modeling
Effect Frame Sampling. To identify the frame that best captures the outcome of a given action, we rank candidate frames within each action segment based on semantic relevance and visual clarity. For semantic relevance, motivated by Niu et al. (2024), we use GPT-4o (Hurst et al., 2024) to generate textual descriptions of the anticipated post-action states. Then we compute the frame-wise similarity between the segment’s visual feature $\mathbf{X}$ and the description embeddings obtained from a pre-trained text encoder, followed by softmax normalization. For visual clarity, we apply the Laplacian operator to estimate the sharpness of each frame within the segment and normalize the scores between the $5th$ and $95th$ percentiles. The final ranking score combines both semantic relevance and visual clarity, and the top-ranked frame is selected as the effect frame.
@@
In the visual branch, we use Grounding DINO (Liu et al., 2024) to detect relevant objects, then we extract object state and relation features $\mathbf{v}_{s},\mathbf{v}_{r}$ by concatenating and projecting embeddings:
@@
Although visual grounding provides object-centric features that capture appearance and spatial cues from the effect frame, it lacks high-level semantic abstraction and relational reasoning. To address this, we introduce a textual scene graph branch to model the symbolic structure of the action effect. Given an effect frame and objects detected by Grounding DINO, we leverage the recognition and reasoning capabilities of a Multimodal Large Language Model (MLLM) by prompting GPT-4o to generate a scene graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ contains three types of nodes: object, relation, and attribute, and $\mathcal{E}$ represents the set of directed edges. We encode the textual labels of all nodes in $\mathcal{V}$ using a pre-trained text encoder to obtain node embeddings, which are then fed into a Graph Neural Network (GNN) (Veličković et al., 2017) to compute context-aware features $\mathbf{t}_{v}$ for each node.
@@ 3.5 Mistake Detection
Following previous work (Lee et al., 2024; Huang et al., 2025a), we adopt an one-class classification (OCC) setting for mistake detection, which learns canonical action patterns from correct samples only. Given the effect-enhanced segment features, we first perform an average-pooling over the enhanced effect-aware segment feature $\widetilde{\mathbf{X}}$ across temporal dimension to obtain the action embedding $\mathbf{x}_{a}$. Instead of learning a binary classifier, we introduce a prompt-based detector to increase task-specific discriminability. For each action label $a$, we construct a template-based textual prompt $\mathcal{P}_{a}$ (e.g., “An image showing [ACTION] for [TASK]”), and obtain its representation by:
@@ 3.6 Training Objective
Finally, we jointly optimize three objectives to train the model in an end-to-end fashion: the segmentation loss $\mathcal{L}^{\text{seg}}$ following Zhang et al. (2022), the action-effect modeling loss $\mathcal{L}^{\text{eff}}$ and $\mathcal{L}^{\text{CL}}$, and the mistake detection loss $\mathcal{L}^{\text{det}}$. The overall training objective is defined as:
@@ 4 Experiments
Table 1: Mistake detection results on the EgoPER dataset (Lee et al., 2024). AUC and EDA are in percentage (%). Best results are bold, and second-best are underlined.
@@
Table 2: Mistake detection results on the CaptainCook4D dataset (Peddi et al., 2023).
@@ 4.1 Experimental Setup
Mistake Datasets. We evaluate our method on two egocentric video datasets: EgoPER (Lee et al., 2024) and CaptainCook4D (Huang et al., 2025a). The EgoPER dataset consists of 385 egocentric cooking videos spanning 5 recipe categories, with a total duration of 28 hours. Among these, 213 videos contain only correct actions, while the remaining 178 include various execution mistakes. Following Lee et al. (2024), we use 80% of the normal videos for training, 10% for validation, and the remaining 10% together with all erroneous videos for testing. The CaptainCook4D dataset contains 384 videos covering 24 recipes, with a total length of 94.5 hours. Among all recordings, 164 are labeled as normal, while the remaining 220 videos include incorrect action steps. Following Huang et al. (2025a), we use all normal recordings for training and the erroneous ones for testing.
Evaluation Metrics. We follow prior work (Lee et al., 2024; Huang et al., 2025a) by using Area Under the Curve (AUC) and Error Detection Accuracy (EDA) to jointly measure model’s ability to distinguish correct and erroneous actions within each video. AUC assesses frame-level discrimination between errors and non-errors, whereas EDA measures segment-level detection accuracy. Both metrics are computed by aggregating detection results across a densely sampled threshold range from 0 to 1, rather than relying on a single fixed threshold, enabling a more comprehensive evaluation of the model’s behavior across the full spectrum of thresholds.
Implementation Details. We use off-the-shelf EVA-02 CLIP (Fang et al., 2024) for all image and text encoding. For the EgoPER dataset, we follow Lee et al. (2024) to use additional features based on Active Object Detection (AOD). For effect frame sampling, the final score is an average of the semantic relevance and visual quality scores. Following Lee et al. (2024) and Huang et al. (2025a), we independently train and test the model in each recipe and report the average of all recipes as the final performance. Both the effect-frame sampling and multimodal knowledge extraction are performed during the data pre-processing stage to avoid additional training overhead. All experiments were conducted with an NVIDIA RTX 6000 Ada GPU (48GB memory). More implementation details can be found in the Appendix A.
@@ 4.2 Comparison with State-of-the-art
We compare our methods with video anomaly detection methods HF2-VAD (Liu et al., 2021), HF2-VAD + SSPCAB (Ristea et al., 2022) and S3R (Wu et al., 2022), as well as mistake detection methods EgoPED (Lee et al., 2024) and AMNAR (Huang et al., 2025a). As shown in Table 1, on the EgoPER dataset, our method outperforms in both metrics on most of the five tasks, with an average improvement of 5.3% on AUC and 2.3% on EDA. The results in Table 2 further demonstrate its effectiveness on the CaptainCook4D dataset, surpassing AMNAR by 2.7% in Precision and 2.3% in AUC. On both datasets, our EDA is occasionally lower than that of AMNAR. This gap is likely due to AMNAR’s use of a dynamic programming block to construct task graphs that mitigate noise in action-segment labels. Since incorporating task graphs lies beyond the scope of our work, we leave it for future exploration.
Table 3: Ablation studies (a–d), action segmentation results (e), and performance comparison with different scene graph generation models (f) on EgoPER dataset (Lee et al., 2024).